# ROTI-GCV: Generalized Cross-Validation for right-ROTationally Invariant Data

### v1

@@ A.2 Useful Results
Isoperimetric inequalities imply that vectors uniformly distributed on $S^{n-1}$ satisfy convex concentration (see [34]) with constant $K/\sqrt{n}$, for some absolute constant $K$, where we note that having sub-Gaussian tails around the median is equivalent to having them around the mean (see e.g. [5]). We can then apply [1, Theorem 2.3] and rescale. ∎
@@
The first equality is clear. The second can be seen to follow from the observation used in the proof of Lemma 1 and the fact that columns of a Haar distributed matrix are uniform on the sphere (see e.g. [30]). Alternatively, recall that for $\mathbf{g}\sim\mathsf{N}(0,\mathbf{I}_{n})$, $\frac{\mathbf{g}}{\|\mathbf{g}\|}\perp\!\!\!\perp\|\mathbf{g}\|$. Hence
@@
We replace $\mathbf{v}_{n}$ with $\frac{\mathbf{g}}{\|\mathbf{g}\|}$ where $\mathbf{g}\sim\mathsf{N}(0,\mathbf{I}_{p}/p)$. Then conditional on $\mathbf{u}_{n}$, we have $\mathbf{g}^{\top}\mathbf{u}_{n}\sim\mathsf{N}(0,\|\mathbf{u}_{n}\|^{2}/p)$. Some standard tail bounds complete this.
$\mathbb{P}\left(|\mathbf{g^{\top}\mathbf{T}_{4}}|>t\mid\mathbf{u}_{n}\right)$ $\leq\sqrt{\frac{2}{\pi}}\frac{\|\mathbf{u}_{n}\|}{t\cdot\sqrt{p}}\exp\left(-t^{2}/2\cdot\frac{p}{\|\mathbf{u}_{n}\|^{2}}\right).$
Since $\limsup_{n\to\infty}\|\mathbf{u}_{n}\|$ is a.s. bounded, the bound above is summable in $n$, so by Borel-Cantelli, the convergence is almost sure. An application of Slutsky finishes, since we know $\|\mathbf{g}\|\xrightarrow{a.s.}1$ by the strong Law of Large Numbers. ∎
@@ A.3 Proof of Theorem 1
$(1-\mathsf{Tr}(\mathbf{S}_{\lambda})/n)^{2}$ $=\left(1-\mathsf{Tr}((\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\mathbf{D}^{\top}\mathbf{D})/n\right)^{2}=\left(\frac{1}{n}\sum_{i=1}^{n}\frac{\lambda}{\lambda+D_{ii}^{2}}\right)^{2}=\gamma^{2}\lambda^{2}v_{\mathbf{D}}(-\lambda)^{2}$
@@
We handle each term. First, $T_{1}$ simplifies easily into a quadratic form which allows us to apply Hanson-Wright (Lemma 7):
@@
This again concentrates around its expectation, which is
$=\frac{\|\boldsymbol{\beta}\|^{2}}{n}\frac{\mathsf{Tr}(\mathbf{D}^{\top}(\mathbf{I}-\mathbf{D}(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\mathbf{D}^{\top})^{2}\mathbf{D})}{p}$
@@
Letting $\mathbf{P}=\mathbf{D}^{\top}(\mathbf{I}-\mathbf{D}(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\mathbf{D}^{\top})^{2}\mathbf{D}$, Hanson-Wright (Lemma 7)implies
@@
We claim that $\limsup_{n\to\infty}\|\mathbf{T}_{4}\|/\sqrt{n}$ is bounded. This follows from the bound above, the law of large numbers applied to $\boldsymbol{\epsilon}$, and the assumption on $\mathbf{D}$. Hence the bound above is summable in $n$ once more, so by Borel-Cantelli, the convergence is almost sure. An application of Slutsky finishes, since we know $\|\mathbf{g}\|\xrightarrow{a.s.}1$ by the strong Law of Large Numbers. The convergence for $T_{3}$ follows directly from [6, Lemma C.3]
@@ A.4 Proof of Theorem 2
$=\frac{n\mathbb{E}[\mathsf{Tr}(\widetilde{\mathbf{D}}^{\top}\widetilde{\mathbf{D}})]}{n^{\prime}p}\left[\frac{1}{n}\boldsymbol{\beta}(\mathbf{I}-(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-1}\mathbf{X}^{\top}\mathbf{X})^{2}\boldsymbol{\beta}\right.$
$\qquad\left.+\frac{1}{n}\boldsymbol{\epsilon}^{\top}\mathbf{X}(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-2}\mathbf{X}^{\top}\boldsymbol{\epsilon}-2\boldsymbol{\epsilon}^{\top}\mathbf{X}(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-2}\mathbf{X}^{\top}\mathbf{X}\boldsymbol{\beta}\right]$
The normalization assumption insures that the prefactor is $1$. The first term corresponds to bias, the second to the variance, and the third is negligible. For the bias,
$B_{\mathbf{X}}(\hat{\boldsymbol{\beta}},\boldsymbol{\beta})$ $=n^{-1}\boldsymbol{\beta}(\mathbf{I}-(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-1}\mathbf{X}^{\top}\mathbf{X})^{2}\boldsymbol{\beta}$
@@
For the third term, we rewrite it as
@@
Lastly, for the variance, we first compute
@@
Then by [6, Lemma C.3], this term satisfies
@@ A.5 Proof of Corollary 4
Combining Equation (15) with Theorem 2 then completes the corollary. ∎
@@ A.7 Proof of Corollary 3
This follows from the fact that the method is continuous in the $t_{i}$ and we have pointwise convergence from Lemma 2.

### v2

@@ A.2 Useful Results
Isoperimetric inequalities imply that vectors uniformly distributed on $S^{n-1}$ have sub-Gaussian tails around their median (see [32]). Explicitly, it is stated that for $\mathbf{v}\sim\mathrm{Unif}(S^{n-1})$, one has that for any $1$-Lispchitz function $f:\mathbb{R}^{n}\to\mathbb{R}$, one has
$\mathbb{P}(|f(\mathbf{v})-M|\geq t)\leq 2e^{-nt^{2}/2},$ (13)
where $M$ denotes the median of $f(X)$ (i.e. $\min(\mathbb{P}(f(\mathbf{v})\geq M),\mathbb{P}(f(\mathbf{v})\leq M))\geq\frac{1}{2}$). We now use [5, Theorem 1.2], which shows that having sub-Gaussian tails with respect to the median is equivalent to having sub-Gaussian tails with respect to the mean, with different constants. Thus, it holds that for any $1$-Lipschitz function $f$, one has
$\mathbb{P}(|f(\mathbf{v})-\mathbb{E}f(\mathbf{v})|\geq t)\leq 4e^{-nt^{2}/32}.$ (14)
Note that this bound is vacuous whenever $14e^{-nt^{2}/32}>$ 1, and hence we can rewrite this as
$\mathbb{P}(|f(\mathbf{v})-\mathbb{E}f(\mathbf{v})|\geq t)\leq\min(4e^{-nt^{2}/32},1)$ (15)
where we now note that for sufficiently large $K$, one has $\min(4e^{-nt^{2}/32},1)\leq 2e^{-nt^{2}/K}$ for any $n\geq 1$ and $t\geq 0$. Hence, one has
$\mathbb{P}(|f(\mathbf{v})-\mathbb{E}f(\mathbf{v})|\geq t)\leq 2e^{-nt^{2}/K}.$ (16)
Finally, [1, Theorem 2.3] implies that random variables satisfying such types of concentration inequalities must in turn satisfy a Hanson-Wright type inequality. That is, Eq. (16) implies that, for fixed matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$, one has
$\mathbb{P}(\left|\mathbf{v}^{\top}\mathbf{A}\mathbf{v}-\mathbb{E}[\mathbf{v}^{\top}\mathbf{A}\mathbf{v}]\right|\geq t)\leq 2\exp\left(-C\min\left(\frac{n^{2}t^{2}}{2K^{4}\|\mathbf{A}\|_{F}^{2}},\frac{nt}{K^{2}\|\mathbf{A}\|}\right)\right)$ (17)
as desired.
@@
The first equality follows from the cyclic property of trace. The second can be seen to follow from the observation used in the proof of Lemma 6 that $\mathbb{E}[\mathbf{o}_{i}\mathbf{o}_{i}^{\top}]=\frac{1}{p}\mathbf{I}_{p}$ and the fact that columns of a Haar distributed matrix are uniform on the sphere (see e.g. [28]). Alternatively, recall that for $\mathbf{g}\sim\mathsf{N}(0,\mathbf{I}_{n})$, $\frac{\mathbf{g}}{\|\mathbf{g}\|}\perp\!\!\!\perp\|\mathbf{g}\|$ from [27, Theorem 2.7.1]. Hence
@@
We replace $\mathbf{v}_{n}$ with $\frac{\mathbf{g}}{\|\mathbf{g}\|}$ where $\mathbf{g}\sim\mathsf{N}(0,\mathbf{I}_{p}/p)$. Then conditional on $\mathbf{u}_{n}$, we have $\mathbf{g}^{\top}\mathbf{u}_{n}\sim\mathsf{N}(0,\|\mathbf{u}_{n}\|^{2}/p)$. We then use Mill’s inequality to show $\mathbf{g}^{\top}\mathbf{u}_{n}\xrightarrow{a.s.}0$:
$\mathbb{P}\left(|\mathbf{g}^{\top}\mathbf{u}_{n}|>t\mid\mathbf{u}_{n}\right)$ $\leq\sqrt{\frac{2}{\pi}}\frac{\|\mathbf{u}_{n}\|}{t\cdot\sqrt{p}}\exp\left(-t^{2}/2\cdot\frac{p}{\|\mathbf{u}_{n}\|^{2}}\right).$
Since $\limsup_{n\to\infty}\|\mathbf{u}_{n}\|$ is a.s. bounded, the bound above is summable in $n$, so by Borel-Cantelli, the convergence is almost sure. An application of Slutsky finishes, since we know $\|\mathbf{g}\|\xrightarrow{a.s.}1$ by the strong Law of Large Numbers, and hence $\mathbf{g}^{\top}\mathbf{u}_{n}/\|\mathbf{g}\|\to 0$. ∎
@@ A.3 Proof of Theorem 1
First recall the statement of Theorem 1: Under the stated assumptions,
$\mathsf{GCV}_{n}(\lambda)-\frac{r^{2}(v_{\mathbf{D}}(-\lambda)-\lambda v^{\prime}_{\mathbf{D}}(-\lambda))+\sigma^{2}\gamma v^{\prime}_{\mathbf{D}}(-\lambda)}{\gamma v_{\mathbf{D}}(-\lambda)^{2}}\xrightarrow{a.s.}0.$
@@
$(1-\mathsf{Tr}(\mathbf{S}_{\lambda})/n)^{2}$ $=\left(1-\mathsf{Tr}((\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\mathbf{D}^{\top}\mathbf{D})/n\right)^{2}=\left(\frac{1}{n}\sum_{i=1}^{n}\frac{\lambda}{\lambda+D_{ii}^{2}}\right)^{2}=\lambda^{2}v_{\mathbf{D}}(-\lambda)^{2}$
@@
We handle each term. First, $T_{1}$ simplifies easily into a quadratic form involving the uniformly random vector $\mathbf{O}\boldsymbol{\beta}$ which is independent of $\mathbf{D}$. This will later allow us to apply Lemma 7:
@@
The expectation of this quantity is as follows:
$\mathbb{E}[T_{1}]$ $=\frac{\|\boldsymbol{\beta}\|^{2}}{n}\frac{\mathsf{Tr}(\mathbf{D}^{\top}(\mathbf{I}-\mathbf{D}(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\mathbf{D}^{\top})^{2}\mathbf{D})}{p}$
@@
Now, letting $\mathbf{P}=\mathbf{D}^{\top}(\mathbf{I}-\mathbf{D}(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\mathbf{D}^{\top})^{2}\mathbf{D}$, Hanson-Wright (Lemma 7) implies
@@
We claim that $\limsup_{n\to\infty}\|\mathbf{T}_{4}\|/\sqrt{n}$ is bounded. This follows from the bound above, the law of large numbers applied to $\boldsymbol{\epsilon}$, and the assumption on $\mathbf{D}$. Hence the bound above is summable in $n$ once more, so by Borel-Cantelli, the convergence is almost sure. An application of Slutsky finishes, since we know $\|\mathbf{g}\|\xrightarrow{a.s.}1$ by the strong Law of Large Numbers. The convergence for $T_{3}$ follows directly from [7, Lemma C.3]
@@ A.4 Proof of Theorem 2
$=\frac{n\mathbb{E}[\mathsf{Tr}(\widetilde{\mathbf{D}}^{\top}\widetilde{\mathbf{D}})]}{n^{\prime}p}\left[\underbrace{\frac{1}{n}\boldsymbol{\beta}(\mathbf{I}-(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-1}\mathbf{X}^{\top}\mathbf{X})^{2}\boldsymbol{\beta}}_{T_{1}}\right.$
$\qquad\left.+\underbrace{\frac{1}{n}\boldsymbol{\epsilon}^{\top}\mathbf{X}(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-2}\mathbf{X}^{\top}\boldsymbol{\epsilon}}_{T_{2}}-\underbrace{2\boldsymbol{\epsilon}^{\top}\mathbf{X}(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-2}\mathbf{X}^{\top}\mathbf{X}\boldsymbol{\beta}}_{T_{3}}\right]$
Remark 8.
The normalization assumption (Assumption A5) ensures that the prefactor is finite. Without further loss of generality, we can assume it is 1. In fact it is sufficient for all proofs to simply assume that lim $\limsup\frac{n\mathbb{E}[\mathsf{Tr}(\widetilde{\mathbf{D}}^{\top}\widetilde{\mathbf{D}})]}{n^{\prime}p}\leq C$, but assuming it is constant simplifies notation.
For intuition why this term is bounded, note that $\mathsf{Tr}(\widetilde{\mathbf{D}}^{\top}\widetilde{\mathbf{D}})=\mathsf{Tr}(\widetilde{\mathbf{X}}^{\top}\widetilde{\mathbf{X}})$, which is the covariance matrix of $n^{\prime}$ samples normalized by $\sqrt{n}$ (the reason for this is given in Remark 2 – since $\boldsymbol{\beta}$ is scaled up by $\sqrt{n}$ all data must be scaled down by $\sqrt{n}$). Hence we expect this trace to be order $n^{\prime}p/n$, which exactly cancels.
The first term corresponds to bias, the second to the variance, and the third is negligible. For the first term,
$T_{1}$ $=n^{-1}\boldsymbol{\beta}(\mathbf{I}-(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-1}\mathbf{X}^{\top}\mathbf{X})^{2}\boldsymbol{\beta}$
@@
For the third term $T_{3}$, we rewrite it as
@@
Lastly, for the second term $T_{2}$, we first compute
@@
Then by [7, Lemma C.3], this term satisfies
@@ A.5 Proof of Corollary 3
Combining Eq. (21) with Theorem 2 then completes the corollary. ∎
@@ A.7 Proof of Corollary 2
A.7.1 Nondegeneracy on the spectrum of $\mathbf{D}$
The exact nondegeneracy condition we require is composed of two portions:
  1. 1.
    Define the events $A_{n}=\bigcup_{i=1}^{L}\{a_{i}=0\}$ and $B_{n}=\bigcup_{i=1}^{L}\{b_{i}=0\}$. Then $\mathds{1}(A_{n}\cup B_{n})\xrightarrow{a.s.}0$.
  2. 2.
    There exists $i_{*}\in\{2,\dots,L\}$ such that lim $\liminf|a_{i_{*}}b_{1}-a_{1}b_{i_{*}}|>C$ almost surely for some $C>0$.
The need for the first condition is clear: our estimator would not even be defined if the entire spectrum were zero; in fact, the problem would be impossible, as the entire data matrix would be zero. The second is more complex, and we give a detailed discussion on its necessity before giving an explicit proof.
Note furthermore that each of these terms can be observed when using the procedure - this gives the user a good way to check if this assumption is likely satisfied in practice.
As motivation, we first discuss the simpler case when $L=2$, as this has connections to the assumptions required for consistent estimation in [23]. In this setting, the estimator for $\sigma^{2}$ becomes
$\frac{\left(\frac{t_{2}}{a_{2}}-\frac{t_{1}}{a_{1}}\right)\left(\frac{b_{2}}{a_{2}}-\frac{b_{1}}{a_{1}}\right)}{\left(\frac{b_{2}}{a_{2}}-\frac{b_{1}}{a_{1}}\right)^{2}}$ $=\frac{\left(\left(\frac{b_{2}}{a_{2}}-\frac{b_{1}}{a_{1}}\right)\sigma^{2}+o(1)\cdot\left(\frac{1}{a_{2}}-\frac{1}{a_{1}}\right)\right)\left(\frac{b_{2}}{a_{2}}-\frac{b_{1}}{a_{1}}\right)}{\left(\frac{b_{2}}{a_{2}}-\frac{b_{1}}{a_{1}}\right)^{2}}$
$=\sigma^{2}+o(1)\frac{a_{1}-a_{2}}{a_{1}b_{2}-b_{1}a_{2}}.$
where the first line comes from Lemma 1. Similarly, the estimator for $r^{2}$ is then $r^{2}+o(1)\frac{b_{2}-b_{1}}{a_{1}b_{2}-b_{1}a_{2}}$. From this and the fact that the $a_{i}$ and $b_{i}$ are all bounded above, we see that this error term is asymptotically negligible provided $a_{1}b_{2}-b_{1}a_{2}$ is bounded away from zero in the limit. In the limiting case where $\lambda_{2}=\infty$, one has $a_{2}=\frac{1}{p}\sum_{i=1}^{p}D_{ii}^{2}$ and $b_{2}=1$, meaning we require
lim $\liminf\left|a_{1}b_{2}-a_{2}b_{1}\right|=\liminf\left|\frac{1}{p}\sum_{i=1}^{p}\frac{\lambda^{2}D_{ii}^{2}}{(D_{ii}^{2}+\lambda)^{2}}-\left(\frac{1}{p}\sum_{i=1}^{p}D_{ii}^{2}\right)\left(\frac{1}{n}\sum_{i=1}^{n}\frac{\lambda^{2}}{(D_{ii}^{2}+\lambda)^{2}}\right)\right|>c>0$ (22)
almost surely for some $c$. This is exactly the second part of our nondegeneracy assumption. Furthermore, this is the exact analog of [23, Assumption 7] in our setting. The above argument thus establishes that consistent estimation is possible if one assumes equation (22).
We do not believe this condition to be simply technical. Consider the setting where $n\leq p$ and every $D_{ii}=1$ for $i\leq n$ (meaning $D_{ii}=0$ for $i>n$). Let $\tilde{\boldsymbol{\beta}}=\mathbf{O}{\boldsymbol{\beta}}$. Then one is given $\mathbf{y}=\mathbf{Q}^{\top}\tilde{\boldsymbol{\beta}}_{[1:n]}+\boldsymbol{\epsilon}$ and the goal is to infer $r^{2}=\|\boldsymbol{\beta}\|_{2}^{2}/\sqrt{n}=\|\tilde{\boldsymbol{\beta}}\|_{2}^{2}/\sqrt{n}$ and $\sigma^{2}$. Every entry $y_{i}$ is now $\mathbf{q}_{i}^{\top}\tilde{\boldsymbol{\beta}}_{[1:n]}+\epsilon_{i}$; the first term has mean zero and variance approximately $r^{2}$, and the second has mean zero and variance $\sigma^{2}$, and it essentially becomes impossible to decouple the effects of $r^{2}$ and $\sigma^{2}$. In this setting, the term above is exactly zero, reflecting that the estimating equations become linearly dependent. That said, we believe this issue should not affect practical use cases; the term we require to be bounded away from zero can be computed by the practitioner directly, and while it is impossible to check if some quantity is bounded away in the limit, they can observe its magnitude.
A.7.2 Proof
Having now overviewed a specific case of our method that yields consistent estimation through the use of an additional assumption already used in the literature, we now turn to the general setting. Note that in the general setting, the condition is weaker – we only require that lim $\liminf|a_{1}b_{i}-a_{i}b_{1}|>C>0$ holds for one of the $i$, meaning one loses nothing by using multiple $\lambda_{i}$, and in fact in practice we observe this results in improved stability.
Here the estimators take on the following forms:
$\hat{\sigma}^{2}$ $=\sigma^{2}+o(1)\frac{\sum_{i=1}^{L}(a_{i}^{-1}-a_{1}^{-1})(b_{i}a_{i}^{-1}-b_{1}a_{1}^{-1})}{\sum_{i=1}^{L}(b_{i}a_{i}^{-1}-b_{1}a_{1}^{-1})^{2}}$
$\hat{r}^{2}$ $=r^{2}+o(1)\frac{\sum_{i=1}^{L}(b_{i}^{-1}-b_{1}^{-1})(a_{i}b_{i}^{-1}-a_{1}b_{1}^{-1})}{\sum_{i=1}^{L}(a_{i}b_{i}^{-1}-a_{1}b_{1}^{-1})^{2}}.$
As a result, we will require an that the coefficient of both error terms is uniformly bounded, which will yield that our estimators are consistent, meaning it suffices to show
lim $\limsup\max\left(\left|\frac{\sum_{i=1}^{L}(a_{i}^{-1}-a_{1}^{-1})(b_{i}a_{i}^{-1}-b_{1}a_{1}^{-1})}{\sum_{i=1}^{L}(b_{i}a_{i}^{-1}-b_{1}a_{1}^{-1})^{2}}\right|,\left|\frac{\sum_{i=1}^{L}(b_{i}^{-1}-b_{1}^{-1})(a_{i}b_{i}^{-1}-a_{1}b_{1}^{-1})}{\sum_{i=1}^{L}(a_{i}b_{i}^{-1}-a_{1}b_{1}^{-1})^{2}}\right|\right)<C$ (23)
almost surely for some $C$.
Now, note that
$a(\lambda)$ $=\frac{\lambda^{2}}{\gamma}\left(v_{\mathbf{D}}(-\lambda)-\lambda v^{\prime}_{\mathbf{D}}(-\lambda)\right)=\frac{1}{p}\sum_{i=1}^{p}D_{ii}^{2}\left(1-\frac{D_{ii}^{2}}{D_{ii}^{2}+\lambda}\right)^{2}$
$b(\lambda)$ $=\lambda^{2}v^{\prime}_{\mathbf{D}}(-\lambda)=\frac{1}{n}\sum_{i=1}^{n}\frac{\lambda^{2}}{(D_{ii}^{2}+\lambda)^{2}}$
are both increasing in $\lambda$, and thus $a_{1}$ is the smallest of the $a_{i}$; likewise for $b_{1}$. Furthermore, note that the $a_{i}$ are bounded above by $\|\mathbf{D}\|_{\mathsf{op}}$ which is then bounded almost surely by assumption; the $b_{i}$ are bounded above by $1$.
We then calculate as follows:
$\frac{\left|\sum_{i=1}^{L}(a_{i}^{-1}-a_{1}^{-1})(b_{i}a_{i}^{-1}-b_{1}a_{1}^{-1})\right|}{\sum_{i=1}^{L}(b_{i}a_{i}^{-1}-b_{1}a_{1}^{-1})^{2}}$ $\leq\frac{\left|\sum_{i=1}^{L}(a_{i}^{-1}-a_{1}^{-1})(b_{i}a_{i}^{-1}-b_{1}a_{1}^{-1})\right|}{(b_{i_{*}}a_{i_{*}}^{-1}-b_{1}a_{1}^{-1})^{2}}\cdot\frac{(a_{1}a_{i_{*}})^{2}}{(a_{1}a_{i_{*}})^{2}}$
$=\frac{\left|\sum_{i=1}^{L}(a_{i_{*}}(a_{1}a_{i}^{-1})-a_{i_{*}})(b_{i}a_{i_{*}}(a_{1}a_{i}^{-1})-b_{1}a_{i_{*}})\right|}{(b_{i}a_{1}-b_{1}a_{i_{*}})^{2}}$
$\leq\frac{1}{(b_{i}a_{1}-b_{1}a_{i_{*}})^{2}}\sum_{i=1}^{L}\left|(a_{i_{*}}(a_{1}a_{i}^{-1})-a_{i_{*}})\right|\left|(b_{i}a_{i_{*}}(a_{1}a_{i}^{-1})-b_{1}a_{i_{*}})\right|$
Note now that since $a_{1}$ is the smallest of the $a_{i}$, $a_{1}a_{i}^{-1}\leq 1$ for all $i$; hence, combined with the fact that the $a_{i}$ and $b_{i}$ are all bounded above in the limit, the final term is also controlled in the limit.
lim $\limsup\frac{\left|\sum_{i=1}^{L}(a_{i}^{-1}-a_{1}^{-1})(b_{i}a_{i}^{-1}-b_{1}a_{1}^{-1})\right|}{\sum_{i=1}^{L}(b_{i}a_{i}^{-1}-b_{1}a_{1}^{-1})^{2}}$ $\stackrel{{\scriptstyle a.s.}}{{\leq}}\frac{1}{C^{\prime}}L\cdot C^{\prime\prime}\cdot 2\cdot C^{\prime\prime}$
where $C^{\prime}$ is the constant from the non-degeneracy assumption and $C^{\prime\prime}$ is the constant from Assumption A2.
The same computation can be performed for the other term in (23) which proceeds exactly as above, but with the $a_{i}$’s and $b_{i}$’s exchanged. The result is
lim $\limsup\frac{\left|\sum_{i=1}^{L}(b_{i}^{-1}-b_{1}^{-1})(a_{i}b_{i}^{-1}-a_{1}b_{1}^{-1})\right|}{\sum_{i=1}^{L}(a_{i}b_{i}^{-1}-a_{1}b_{1}^{-1})^{2}}$ $\stackrel{{\scriptstyle a.s.}}{{\leq}}\frac{1}{C^{\prime}}L\cdot 2\cdot C^{\prime\prime},$
as desired.